Two boys are holding a horizontal rod of length L and weight W through its two ends. If now one of the boys suddenly leaves the rod what is the instantaneous reaction force experienced by the other boy?

Grade:12th pass

1 Answers

Sakshi
6 years ago
At centre of mass torque will beR.l/2=ML*2/12@About origin torque will beW.l/2=ML*2/3@Divide 1 $ 2R=W/4